Calculate the momentum change for the box with mass m = 500 g during the delta time t = 0.5 s, when it is known that the coefficient of sliding friction is u = 0.1.


1
Expert's answer
2021-01-27T09:30:44-0500

If only the force of friction acts on the body then


"F=\frac{\Delta p}{\Delta t}\to \Delta p=F\cdot\Delta t="


"=\mu mgt=0.1\cdot0.5\cdot9.8\cdot 0.5=0.245(kg\cdot m/s)" . Answer